A ball is launched from the lab table, with an initial velocity of 2.5 m/s. If the table is 0.7 m high, how far from the table does it land? How long does it take to land?
1
Expert's answer
2020-12-14T07:26:42-0500

If "\alpha=0" then


"v_{0x}=2.5 (m\/s)" and "v_{0y}=0"


"h=gt^2/2\to"


"t=\sqrt{2h/g}=\sqrt{2\cdot 0.7/9.8}=0.34(s)" .Answer


"x=v_{0x}t=2.5\cdot0.34=0.85(m)". Answer